OpenStudy (anonymous):

suppose y varies directly with x, and y=10 when x= -3 what direct variation equation relates x and y? what is the value of y when x= -1?

OpenStudy (hba):

\[\huge\ y \alpha x\] \[\huge\ y=kx\] \[\huge\ k=y/x\] Here y=10 and x=-3 so just plug them in :)

OpenStudy (hba):

So what do you get :)

OpenStudy (hba):

Where are you confused ?

OpenStudy (hba):

Well let me explain y varies directly with x means \[y \alpha x\] Where \[\\ \alpha \ is \\ the \\ sign \\ of \\ proportionality.\] So when we remove this sign we have to put a constant with the equals sign So let the constant is k so, \[y=kx\] and \[k=y/x\]